# Not All Frame Features Are Equal: Video-to-4D Generation via Decoupling Dynamic-Static Features

### v1

@@ 3.5 Training Objectives
Following [1], we calculate SDS loss based on rendered views and ground truth images, including input video and pseudo multi-view images. Furthermore, we compute the photometric loss among them. Considering if we employ photometric loss on the pseudo multi-view images with rendered views, the quality of pseudo multi-view images directly impacts optimization on pixel-level. Therefore, we introduce the LPIPS loss [54] widely used in 3D generation [55, 56, 19] to minimize the similarity between pseudo multi-view images and rendered views. More details can be found in the supplementary material.
@@ 4.1 Implementation Details
Datasets For a fair quantitative evaluation, we utilize the dataset provided by Consistent4D. Consistent4D dataset includes 7 dynamic objects. Each object contains an input video and four ground truth videos with different camera settings. Additionally, we introduce an extra 7 dynamic objects sampling from a subset of Objaverse [21, 13]. Each object is rendered to 24 input frames with a front camera pose and 24 ground truth frames with 360 cameras. Furthermore, we use a set of videos from online sources for qualitative evaluation. In Sec.5, to evaluate the effectiveness of our method in real-world scenarios, we utilize three real-world scenarios provided by Neu3D’s [2] dataset.
@@
Training Our two models, including DS4D-GA (using GA in TSSF) and DS4D-DA (using DA in TSSF) use the same training setting. During the initial 1,000 iterations, we train our models except TSSF and deformation MLP. Subsequently, the models with TSSF and deformation MLP are optimized over 6,000 additional iterations. For the deformation MLP, we employ each MLP with 64 hidden layers and 32 hidden features. The learning rate of TSSF and deformation MLP is set to $1.6\times 10^{-4}$ and is decayed to $1.6\times 10^{-6}$. Following [1], the top $2.5\%$ of points are densified with the most accumulated gradient. The overall training process costs approximately 3 hours on a V100 GPU.
@@ 4.2 Comparisons with Existing Methods
Quantitative comparisons. The quantitative results are shown in Tab. 1. Our methods DS4D-GA and DS4D-DA, consistently outperform other methods in all metrics. Specifically, our methods notably exceed the SOTAs in FVD, indicating that our generation results have fewer temporal artifacts than others. Furthermore, the experiments on Objaverse dataset present that our methods significantly outperform other SOTAs by a large margin. It demonstrates our methods are superior in terms of temporal-spatial consistency, fidelity, and quality of generation results and underscores the robustness of our methods. In summary, such significant improvements are attributed to our method of decoupling dynamic-static features, which can explicitly distinguish dynamic and static zones within frame features.
Qualitative comparisons. The qualitative results on Consistent4D dataset are presented in Fig. 5. Besides, the results on Objaverse dataset are shown in Fig.6. Obviously, whether in the example with a larger proportion of static regions (the person in Fig. 5 and warrior in Fig.6) or the example with a balanced proportion of dynamic and static regions (guppie in Fig. 5), the results, generated by Consistent4D, Dreamgaussian4D, STAG4D and SC4D, have different degrees of blurriness in the details, especially in the areas with motion trends in the current and subsequent frames. Although 4Diffusion and L4GM generate videos with clearer textures than other previous methods with the help of a large amount of 4D data prior, it is easy to generate the abnormal shape or textures with inconsistent details in some areas (e.g, the back of the person). These methods do not differentiate between dynamic and static information, leading them to easily overlook information in dynamic regions when faced with large proportions of static. In contrast, our methods address the challenges arising from varying proportions of dynamic and static regions in frames, leading to the achievement of high-quality 4D generation.
@@ 4.3 Ablation Experiments
Figure 8: (a) Visualization on the heatmap of dynamic features in DSFD. The red region highlights the primary zone of interest in dynamic features. (b) The score map of point features in TSSF-GA and TSSF-DA. The red area indicates high attention to dynamic information in point features of a specific view.
Effect of DSFD. To validate the effect of DSFD, we conduct experiments in two aspects: a) Adding frame features in model (labeled as D). b) Performing decoupling on frame features in model (labeled as E, also termed DSFD). In particular, the performance of D and E both have significant improvement compared to A and B, especially FVD score. For D, the frame features provide enough structure prior knowledge and texture information for Deformation MLP. It assists Deformation MLP in capturing shape deformation. Nevertheless, D fails to differentiate between dynamic and static regions of frames clearly. Consequently, D is prone to overfitting in static areas, resulting in blurry details. For example, the blurred details of legs from D in Fig.7. In contrast, DSFD which decouples dynamic-static features, mitigates the issue.
Effect of TSSF. To validate the effect of TSSF, we add the TSSF based on model E, including model F, G (TSSF-GA) and H (TSSF-DA). Among them, we replace adaptive fusion with average pooling in model F. Specifically, we observe that E, G, H can generate more texture details than other models. Because they mitigate the problem brought by spatial occlusion by aggregating dynamic components from different viewpoints. However, F only averages the point features from DSFD along spatial axes, which ignores merging similar dynamic information. The model G and F adaptively select dynamic information, which plays a crucial role in enhancing dynamic representations. Compared with E, the texture of 4D content generated by F and H is more refined in Fig.7. This underscores the efficacy of TSSF, including TSSF-GA and TSSF-DA.
@@ 4.4 Visualization
Decoupled in DSFD. Fig.8 (a) presents the heatmap of dynamic features obtained by DSFD decoupling features from the current and reference frame features. The red area indicates the primary region of interest in the features. We observe the motion trends come from the trunk and body of elephant in the front view (left of the figure). Thus, dynamic features show strong concern in this area. Limited to the viewpoint range, the motion trajectory for pseudo views (right of the figure) is more interested in the elephant’s head. No doubt, dynamic features also present a high response in similar areas. In conclusion, our method can acquire accurate dynamic features using DSFD, as supported by the visualization results.
Selected in TSSF. Fig.8 (b) shows the score map of selecting similarity dynamic information from point features of different views by TSSF-GA and TSSF-DA. Specifically, the leg movements of the triceratops indicate the primary trend in motion between the middle and current time. The red area indicates the high attention to dynamic information in point features of a specific view. Meanwhile, this area is also similar to the dynamic area of other views in Fig.8 (b). Hence, it reveals that the two approaches can capture a certain degree of similar dynamic information from different viewpoints. However, TSSF-GA is interested in the back of triceratops rather than legs since the front legs are nearly invisible in the pseudo views $v_{2}$. The unseen dynamic region in pseudo view influences the accuracy of selecting similar dynamic information. In contrast, TSSF-DA predicts highest scores on the legs under view $v_{1}$ and $v_{2}$, thanks to reducing the impact of pseudo views in TSSF-DA. It indicates TSSF-DA can alleviate issues caused by pseudo views.
@@ 5 Discussion on Real-World Scenario
Generating a dynamic scene from input real-world videos is an essential task. In this section, we conduct experiments on three scenes from Neu3D’s dataset to demonstrate the effectiveness of our proposed on real-world dynamic scene. We set 4D-GS [9] as the baseline, which has shown promising results in this task. To fair comparison, we directly insert our DSFD and TSSF into 4D-GS. More details can be found in supplementary material.
@@
Figure 9: Visualization of flame steak compared with 4D-GS.
@@
6 Conclusion and Limitations
In this paper, we propose a novel framework DS4D, which decouples dynamic-static information along temporal-spatial axes to enhance dynamic representations for high-quality 4D generation. Its DSFD decouples features by regarding the portions of current features that possess significant differences relative to reference frame features as dynamic features. Moreover, TSSF is designed to enhance dynamic representations by selecting similar dynamic information. Overall, our method can produce high-quality 4D content and shows promise in 4D scene generation.
Limitations. There are sevral limitations in our method. 1) Since our DS4D relies on multi-view diffusion, e.g., zero123++ [18], the quality of pseduo multi-view squences still influence our generation results (More experiments and analysis can be found in supplementary material). 2) Limited to the resolution of input video, it is challenge for our method to produce high-resolution 4D contents , e.g., over 2K resolution.

### v2

@@ 3.5 Training Objectives
Following [1], we calculate SDS loss and the photometric loss based on rendered views and ground truth images. Inspired by [hong2023lrm, zou2024triplane, 19], we introduce the LPIPS loss to minimize the similarity between pseudo multi-view images and rendered views. More details can be found in the supplementary material.
@@ 4 Experiments
Figure 5: Qualitative comparison on video-to-4D generation. For each method, we render results under two novel views at two timestamps. More comparison examples, and some of the 360 generation results (video file) can be found in supplementary materials.
@@ 4.1 Implementation Details
Datasets In our article, we use four challenging datasets. 1) We utilize the dataset provided by Consistent4D. 2) We introduce a subset of Objaverse [21, 13]. The objects in Objaverse have more complexity motion than Consistent4D dataset. 3) We introduce some challenging videos from online sources. 4) In Sec.5, to evaluate the effectiveness of our method in real-world scenarios, we utilize three real-world scenarios provided by Neu3D’s [2] dataset. More details can be found in supplementary material.
@@
Training Our two models, including DS4D-GA (using GA in TSSF) and DS4D-DA (using DA in TSSF) use the same training setting. The training details can be found in Supplementary Material.
@@ 4.2 Comparisons with Existing Methods
Quantitative comparisons. The quantitative results are shown in Tab. 1. Our methods DS4D-GA and DS4D-DA, consistently outperform other methods in all metrics. Specifically, our methods notably exceed the SOTAs in FVD, indicating that our generation results have fewer temporal artifacts than others. Furthermore, the experiments on Objaverse dataset present that our methods significantly outperform other SOTAs by a large margin. It demonstrates our methods are superior in terms of temporal-spatial consistency, fidelity, and quality of generation results and underscores the robustness of our methods. In summary, such significant improvements are attributed to our method of decoupling dynamic-static features, which can explicitly distinguish dynamic and static regions within frame features.
Qualitative comparisons. The qualitative results on Consistent4D dataset are presented in Fig. 16. Besides, the results on Objaverse dataset are shown in Fig.6. Obviously, whether in the example with a larger proportion of static regions (the person in Fig. 16 and warrior in Fig.6) or the example with a balanced proportion of dynamic and static regions (guppie in Fig. 16), the results, generated by Consistent4D, Dreamgaussian4D, STAG4D and SC4D, have different degrees of blurriness in the details, especially in the areas with motion trends in the current and subsequent frames. Although 4Diffusion and L4GM generate videos with clearer textures than other previous methods with the help of a large amount of 4D data prior, it is easy to generate the abnormal shape or textures with inconsistent details in some areas (e.g, the back of the person). These methods do not differentiate between dynamic and static information, leading them to easily overlook information in dynamic regions when faced with large proportions of static. In contrast, our methods address the challenges arising from varying proportions of dynamic and static regions in frames, resulting in the achievement of high-quality 4D generation.
@@ 4.3 Ablation Experiments
Figure 8: (a) Visualization on the heatmap of dynamic features in DSFD. The red region highlights the primary region of interest in dynamic features. (b) The score map of point features in TSSF-GA and TSSF-DA. The red area indicates high attention to dynamic information in point features of a specific view.
Effect of DSFD and Decoupling. Validating the effect of DSFD also means Validating the effect of decoupling since DSFD consists of decoupling. Therefore, we conduct experiments in two aspects: a) Adding frame features in the model without decoupling (labeled as D). b) Performing decoupling on frame features in the model (labeled as E, also termed DSFD). In particular, the performance of D and E both have significant improvement compared to A and B, especially FVD score. For D, the frame features provide enough structure prior knowledge and texture information for Deformation MLP. It assists Deformation MLP in capturing shape deformation. Nevertheless, D fails to differentiate between dynamic and static regions of frames clearly. Consequently, D is prone to overfitting in static areas, resulting in blurry details. For example, the blurred details of legs from D in Fig.7. In contrast, DSFD, which decouples dynamic-static features, mitigates the issue.
Effect of TSSF. To validate the effect of TSSF, we add the TSSF based on model E, including model F, G (TSSF-GA) and H (TSSF-DA). Among them, we replace adaptive fusion with average pooling in model F. Specifically, we observe that E, G, H can generate more texture details than other models. Because they mitigate the problem brought by spatial occlusion by aggregating dynamic components from different viewpoints. However, F only averages the point features from DSFD along spatial axes, which ignores merging similar dynamic information. The model G and F adaptively select dynamic information, which plays a crucial role in enhancing dynamic representations. Compared with E, the texture of 4D content generated by F and H is more refined in Fig.7. This underscores the efficacy of merging similar dynamic information in TSSF, including TSSF-GA and TSSF-DA.
@@ 4.4 Visualization
DSFD. Fig.8 (a) presents the heatmap of dynamic features obtained by DSFD decoupling features from the current and reference frame features. The red area indicates the primary region of interest in the features. We observe the motion trends come from the trunk and body of elephant in the front view (left of the figure). Thus, dynamic features show strong concern in this area. Limited to the viewpoint range, the motion trajectory for novel views (right of the figure) is more interested in the elephant’s head. No doubt, dynamic features also present a high response in similar areas. In conclusion, our method can acquire accurate dynamic features using DSFD, as supported by the visualization results.
TSSF. Fig.8 (b) shows the score map of adaptively selecting similarity dynamic information from point features of different views by TSSF-GA and TSSF-DA. Specifically, the leg movements of the triceratops indicate the primary trend in motion between the middle and current time. The red area indicates the high attention to dynamic information in point features of a specific view. Meanwhile, this area is also similar to the dynamic area of other views in Fig.8 (b). Hence, it reveals that the two approaches can capture a certain degree of similar dynamic information from different viewpoints. However, TSSF-GA is interested in the back of triceratops rather than legs since the front legs are nearly invisible in the novel view $v_{2}$. The unseen dynamic region in the novel view influences the accuracy of selecting similar dynamic information. In contrast, TSSF-DA predicts highest scores on the legs under view $v_{1}$ and $v_{2}$, thanks to reducing the impact of novel views in TSSF-DA. It indicates TSSF-DA can alleviate issues caused by the fact that dynamic regions are noticeably obscured in novel views.
Figure 9: Visualization of flame steak compared with 4D-GS.
@@ 5 Discussion on Real-World Scenario
Generating a dynamic scene from input real-world videos is an essential task. In this section, we conduct experiments on three scenes from Neu3D’s dataset to demonstrate the effectiveness of DSFD and TSSF on complexity real-world dynamic scene. We set 4D-GS [9] as the baseline, which has shown promising results in this task. To fair comparison, we directly insert our DSFD and TSSF into 4D-GS. More details can be found in supplementary material.
@@
6 Conclusion
In this paper, we propose a novel framework DS4D, which decouples dynamic-static information along temporal-spatial axes to enhance dynamic representations for high-quality 4D generation. Its DSFD decouples features by regarding the regions of current features that possess significant differences relative to reference frame features as dynamic features. Moreover, TSSF is designed to enhance dynamic representations by selecting similar dynamic information. Overall, our method can produce high-quality 4D content and shows promise in 4D scene generation.
@@ 6 Conclusion
Supplementary Material
7 The 360 Results Generated by Our methods
We strongly recommend checking the video file in the supplementary material. Some of the 360 results have been attached in the video of the supplementary material. In future official versions, we will provide more 360 results and comparisons.
8 Training Objectives
In this section, we present more details on training objectives. Please note that each loss in our loss function has a clear derivation from previous works. The theoretical derivation can be found in corresponding previous works, e.g., [53, 54, 1] etc. Hence, we will not provide a detailed derivation process in our manuscript.
Score distillation sampling loss. Following [1], we employ multi-view score distillation sampling (SDS) loss using rendered images under camera poses of pseudo multi-view images and input video.In detail, there are rendered multi-view images $I=\{I^{(i,1)},\dots,I^{(i,j)}\}$ under 6 camera poses, where $i$ denotes the timestamps and $j$ denotes the number of views. Formally, SDS loss can be defined as:
$\mathcal{L}_{SDS}$ $=\alpha_{1}\mathcal{L}_{SDS}^{pseudo}+\alpha_{2}\mathcal{L}_{SDS}^{real}$ (6)
$=\alpha_{1}\mathcal{L}_{SDS}(\phi,I^{(i,j)})+\alpha_{2}\mathcal{L}_{SDS}(\phi,I_{real}^{i})$
where $\alpha_{1}$ and $\alpha_{2}$ are hyperparameters, $I_{real}^{i}$ is the rendered image under camera pose of input video at time $i$.
Photometric loss. Following [54, 1], we compute the reconstruction loss $\mathcal{L}_{rec}$ between rendered images and pseudo multi-view images, and the foreground mask $\mathcal{L}_{mask}$.
LPIPS loss. We introduce the LPIPS loss $\mathcal{L}_{lpips}$ [53] to compute the feature similarity between pseudo multi-view images and corresponding rendered images. We leverage VGG [55] as backbone to extract image features.
Overall loss. Based on above loss functions, we obtain the final overall loss $\mathcal{L}$:
$\mathcal{L}=\lambda_{1}\mathcal{L}_{SDS}+\lambda_{2}\mathcal{L}_{rec}+\lambda_{3}\mathcal{L}_{mask}+\lambda_{4}\mathcal{L}_{lpips}$ (7)
where $\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}$ are hyperparameters.
9 Training Details of DS4D-GA and DS4D-DA
We train our two models DS4D-GA (using GA in TSSF) and DS4D-DA (using DA in TSSF) under the same training setting. During the initial 1,000 iterations, we train our models except TSSF and deformation MLP. Subsequently, the models with TSSF and deformation MLP are optimized over 6,000 additional iterations. For the deformation MLP, we employ each MLP with 64 hidden layers and 32 hidden features. The learning rate of TSSF and deformation MLP is set to $1.6\times 10^{-4}$ and is decayed to $1.6\times 10^{-6}$. Following [1], the top $2.5\%$ of points are densified with the most accumulated gradient. The overall training process costs approximately 3 hours on a V100 GPU.
Figure 10: Example on low-quality images generated by ImageDream and high-quality images generated by Zero123++, both using the same input video.
  Methods High-quality Input Images Low-quality Input Images
STAG4D [1] 0.9078 0.1354 986.8271 26.3705 0.9026 0.1437 1311.8770 40.8664
DS4D-GA (Ours) 0.9206 0.1311 799.9367 26.1794 0.9195 0.1380 849.8154 25.8726
DS4D-DA (Ours) 0.9225 0.1309 784.0235 24.0492 0.9221 0.1339 805.4721 24.0623
Table 4: Evaluation and comparison of the performance when facing low-quality input images and high-quality input images. The best score is highlighted in bold.
10 Training Details on Discussion
In this section, we present the training details of experiments about Discussion (Section 5) in our manuscript. For a fair comparison, the training settings are the same as 4D-GS [9].
Figure 11: Qualitative comparison on 4D generation results with low-quality inputs. For each method, we render results under two novel views at three timestamps.
Figure 12: Comparison regarding whether using point clouds generated by a large reconstruction model as initialization. A: The model without point initialization. B: The model with point initialization.
Network Architecture. We add our proposed DSFD in 4D-GS. Meanwhile, we directly replace the spatial-temporal structure encoder and multi-head Gaussian deformation decoder of 4D-GS with our TSSF. Since the multi-view sequences from Neu3D’s dataset are authentic, we use TSSF-GA instead of TSSF-DA. Additionally, the hyperparameter settings of HexPlane are the same as 4D-GS. In detail, the basic resolution of HexPlane is 64, and is unsampled by 2 and 4.
Training Settings. We use the same dense point cloud generated by SFM as 4D-GS. Then we also downsample point clouds lower than 100k. The learning rate of TSSF and deformation MLP are set to $1.6\times 10^{-3}$ which decreases to $1.6\times 10^{-5}$. The batch size is 1. Following 4D-GS, we do not use opacity reset operation. The overall training iterations spend 14000 for each scene.
Note that: a) Our DS4D includes point initialization (init.) via LRM. However, in Neu3D’s data, common point cloud init. (e.g., 4D-GS) uses colmap rather than LRM. Thus, it is unfair to compare DS4D and 4D-GS due to the different init.. b) DSFD and TSSF are core contributions, thus inserting them into 4D-GS can directly demonstrate the effectiveness of our contributions in real-world scenes.
11 More details on Datasets
In this section, we present more details on four datasets.
Consistent4D Dataset. Following [1], we use seven 30-frame video in front view as input video, and their ground-truth with four novel views (azimuth angles of $-75^{\circ}$, $15^{\circ}$, $105^{\circ}$, and $195^{\circ}$, respectively) as evaluation.
Objaverse Dataset. We random sample seven dynamic objects from [21, 13]. The 24-frame ground truth under 360 cameras (the range of azimuth angles is $[0^{\circ},360^{\circ}]$) rendered from each object is used as evaluation. Meanwhile, we use seven 24-frame videos in front view as input video. Compared to Consistent4D Dataset, the object in Objaverse Dataset has more complex motion, e.g., suddenly waving at some time.
Neu3D’s Dataset. We use three real-world scenarios from Neu3D’s dataset [2]. Each scene has 300 frames with 20 cameras, a total of 6000 high-quality images. Following [9], we use 300 frames under the front camera pose as evaluation, others are used as training videos.
Data from Online Sources. Following [1], we introduce some challenging videos from online sources for qualitative evaluation. Moreover, we also generate input videos by Stable Video Diffusion [56]. Each input video has 14 or 30 frames.
12 Experiments
In this section, we conduct more experiments to evaluate our method.
12.1 Robustness of Our Methods
In this section, we explore whether the quality of input images has a huge influence on the generation results of our methods.
Specifically, we construct a dataset with low-quality input multi-view images. Using the same input video as Tab.1 of the main manuscript, we leverage ImageDream [57] to produce a series of multi-view sequences. Then, we select multi-view images with inconsistency or shape deformation from the generation multi-view sequences. These low-quality multi-view images are grouped as the low-quality input images. The example data of low-quality inputs can be seen in Fig. 10. The low-quality inputs has serious inconsistency between different timestamps and has color fading and texture blurry compared with the high-quality inputs we used in Tab.1 of the main manuscript (e.g., the shape and color of the suit).
Based on the low-quality inputs, we compare our methods with STAG4D. The quantitative and qualitative results are shown in Tab. 4 and Fig. 11. Our methods when using low-quality inputs maintain a similar performance compared to our results when using high-quality inputs. However, in image metrics (LPIPS) and video metrics (FID and FID-VID), STAG4D when using low-quality inputs has a significantly worse performance compared to STAG4D using high-quality inputs. Furthermore, in Fig. 11, STAG4D generates the blurry textures in the back view. The reason is that back’s texture details are blurry in the low-quality input at some timestamps (e.g., at $t_{1}$ of $v_{3}$). In contrast, our methods can generate results with clear textures. This indicates that our methods can handle the low-quality inputs better than STAG4D. Since our methods decouple dynamic-static information at the feature-level. Even though input low-quality data, thanks to the robustness feature extraction ability of DINOv2, we can still leverage the inherent differences between features to decouple. The differences include the change of motion, shape and textures between the corresponding two frames.
In summary, the above experiments demonstrate that the quality of input images has few influence on the generation results of our methods and verify the robustness of our methods when input low-quality data.
12.2 More Analysis on Ablation Experiments
Figure 13: Visualization on the heatmap of dynamic features in DSFD. The red region highlights the primary zone of interest in dynamic features.
Figure 14: Visualization on the score map of point features in TSSF-GA and TSSF-DA. The red area indicates model’s high attention on dynamic information in point features of a specific view.
The effect on points initialization. To validate the effect of point initialization, we add the point initialization to the baseline model, which is labeled as B (as shown in Table.2 of the manuscript). After performing point initialization, the performance of B model has improved on all metrics. Additionally, we visualize the Gaussian points from baseline model (labeled as A, as shown in Table.2 of the manuscript) and B in Fig.12. Obviously, the Gaussian points of B are relatively uniform and denser in distribution than those of A, which ensures the stability of optimization and fidelity of the motion and shape fully learned by the model.
The effect on LPIPS Loss. To validate the effect of LPIPS Loss, we add the LPIPS Loss based on model B, labeled as C (as shown in Table.2 of the manuscript). The performance of C model has improved on all metrics. It indicates the effectiveness of LPIPS loss.
12.3 More Visualization
In this section, we provide more visualizations of features in DSFD and TSSF, respectively.
DSFD. In Fig.13, we supplement more heatmaps of dynamic features obtained by DSFD decoupling features from the current and reference frame features. The red area indicates the primary region of interest in the features. The dog’s head movements indicate the main motion trends between the reference frame and the current frame. No matter what kind of novel views, dynamic features decoupled by our DSFD can accurately represent the dynamic zones. It once again demonstrates our method can acquire accurate dynamic features using DSFD.
TSSF. In Fig.14, we supplement more score maps of selecting similarity dynamic information from point features of different views by TSSF-GA and TSSF-DA. Moreover, Fig.15 shows the corresponding generation results at the current timestamp, including the same example as Figure.8 (b) in the manuscript and the same example in Fig.14.
Specifically, the head movements of the person indicate the primary trend in motion between the middle and current time. The red area indicates the model’s high attention on dynamic information in point features of a specific view. According to score maps based on different views, two approaches can capture a certain degree of similar dynamic information from different viewpoints. TSSF-GA is interested in both body and head in $v_{1}$, but TSSF-DA pays more concentration to the head. This is because TSSF-DA reduces the impact of novel views, resulting in TSSF-DA focusing more on capturing regions with a motion trend that is more similar to the front frame in other novel views. Thus, compared to TSSF-GA, TSSF-DA can produce results with clear details in the corresponding regions. For example, the hair texture of person at the top of Fig.15, and the leg texture of triceratops at the down of Fig.15. It once again indicates TSSF-DA can alleviate issues caused by novel views.
Figure 15: Qualitative comparison between model using TSSF-GA and model using TSSF-DA at current timestamp. Their corresponding visualization on the score map of point features is as shown in Fig.14 and Figure.8 (b) in our manuscript.
Figure 16: More Results for 4D Generation using DS4D-GA and DS4D-DA.
(a) Sear steak
(b) Cook spinach
Figure 18: Visualization of real-world 4D scene generation compared with 4D-GS.
12.4 More Qualitative Comparisons
We supplement more qualitative comparisons in Fig.17. We compare our methods DS4D-GA and DS4D-DA with other SOTA methods, including Consistent4D [20], Dreamgaussian4D [47], STAG4D [1], SC4D [16], 4Diffusion [11], and L4GM [12]. All the experiments of the methods are carried out using the code from their official GitHub repository.
12.5 More Examples on 4D Content Generation
We supplement more 4D content generation examples produced by DS4D-GA and DS4D-DA in Fig.16.
12.6 More Examples on Real-World Scenario
We supplement more real-world 4D scene generation examples using our method and baseline 4D-GS [46] in Fig.18. The experiments of 4D-GS are carried out using the code from their official GitHub repository.
12.7 Time Consuming on Decoupling
As mentioned in our manuscript, direct decoupling always costs considerable computation time. To intuitively evaluate the time-consuming advantage of our decoupling approach compared to direct decoupling, we provide the running time of each approach on decoupling dynamic-static features from a frame features with a 30-frame video, as shown in Tab.5. Undoubtedly, our decoupling approach is about 14 times faster than direct decoupling.
Direct Decoupling Our Decoupling Approach
118.326 (ms) 8.241 (ms)
Table 5: Comparison of running time with different decoupling approaches. All approaches are tested on a NVIDIA 3090 GPU.
13 Limitations
Limited to the resolution of input video, it is challenge for our method to produce high-resolution 4D contents , e.g., over 2K resolution.